OpenStudy (anonymous):

Find the vertex and axis of the symmetry of the graph of the function. f(x) =x^2-8x

OpenStudy (shamil98):

Graph the equation, the turning point of the graph is the vertex, the x coordinate of the vertex is the axis of symmetry.

OpenStudy (anonymous):

How do I find the vertex and x coordinate?

OpenStudy (shamil98):

You graph the equation

OpenStudy (anonymous):

No, there is more to that. It's a question on my homework and there isn't a graph

OpenStudy (shamil98):

This is what the graph is.

OpenStudy (shamil98):

The axis of symmetry is x = 4 , and the vertex is (4,-16)

OpenStudy (anonymous):

how did you find that?

OpenStudy (shamil98):

I graphed the equation, f(x) = x^2 -8x just means y = x^2 - 8x

OpenStudy (anonymous):

You could also put it into the completed square form and it will give you the axis of symmetry which you can use to find the vertex

OpenStudy (anonymous):

hmmm I still not sure how to do that. This qustion on my homework doesn't have a graph. How do you do it without a graph?

OpenStudy (shamil98):

If your assignment asks you to find the vertex and axis of symmetry without a graph of it, i assume your class has taught you how to graph functions?

OpenStudy (anonymous):

yes

OpenStudy (shamil98):

writing the equation in y = a(x – h)2 + k gives the vertex, (h,k) are the coordinates to the vertex.

OpenStudy (shamil98):

the standard equation of a parabola is y = ax2 + bx + c.

OpenStudy (anonymous):

Can you please show me the steps on how you came up with the answer without the graph please?

OpenStudy (shamil98):

Okay.

OpenStudy (shamil98):

y = x^2 -8x To find the x coordinate of the vertex ( the axis of symmetry )we use the equation -b/2a. so -(-8)/2(1) 8/2 = 4

OpenStudy (shamil98):

x = 4.

OpenStudy (shamil98):

then substitute that value into y = x^2 - 8x to get the y-coordinate of the vertex y = (4)^2 - 8(4) y = 16 - 32 y= -16

OpenStudy (shamil98):

does this make sense?

OpenStudy (anonymous):

Thank you so much. It's easier for me to understand when I see it.